OpenStudy (anonymous):

Simplify. Write the answer in positive exponents. (15x^-7 y^5)/(5xy^-4)

OpenStudy (anonymous):

(15x^-7 y^5)/(5xy^-4) (15 y^5)* y^4/(5x)x^7 (15 y^9)*/(5x^8)

OpenStudy (ash2326):

we have \[\frac{15x^{-7} y^5}{5xy^{-4}}\] let's bring x^(-7) to the denominator and y^(-4) to the numerator we get\[\frac{15y^5\times y^4}{5x^7\times x}\] we get \[3 \frac{y^{5+4}}{x^{7+1}}=3\frac{y^9}{x^8}\]